Standard for guest rooms:   triple sheet all beds.

Room type: King suite (king bed only)

Number of King suites:    15

Par level:                            4.5

How many king sheets are required for 4.5 par?
Answer #1

Number of sheets per bed = 3

Number of king suits= 15

Total number of sheets per par number= 15*3= 45

Total number of king sheets required for 4.5 par = 45*4.5 = 202.5

Hence 203 king sheets are required for 4.5 par.